OpenStudy (anonymous):

A solution is made by dissolving 15.5 grams of glucose (C6H12O6) in 245 grams of water. What is the freezing-point depression of the solvent if the freezing point constant is -1.86 °C/m? Show all of the work needed to solve this problem

OpenStudy (matt101):

The equation for freezing point depression is: \[\Delta T_f=iK_fm\] Where ΔT(f) is the change in the freezing point, i is the van't Hoff factor, K(f) is the freezing point constant, and is the molality of the solution. Do you think you can solve this question now?

OpenStudy (anonymous):

I think so... Um, if you're finding the molality, you do m of solute/kg of solvent. .17/.245=.69 Then you just plug everything in and multiply it. \[-1.86\times.69=-1.28\] ???

OpenStudy (anonymous):

Don't you have to add the two now? Because when googled it, someone said to add them at the end.

OpenStudy (anonymous):

@matt101

OpenStudy (aaronq):

the formula is \(\sf molality =\dfrac{moles~of~solute}{kg~of ~solution}\) not mass. you need to convert the mass of solute to moles first.

OpenStudy (anonymous):

okay, my bad, so the amount of moles is .09 i'm pretty sure. then when you divide it by .245 it's equal to .37. And then you would plug in a multiply all of them. So \[Delta T =-1.86\times.37\] \[\Delta T =-.69\]

OpenStudy (anonymous):

@aaronq

OpenStudy (matt101):

That's correct!